y = 2x + 7
ordered pairs(0, )
and
(3, )

Answer by nerdybill(7384)

Ordered pairs are given with the x-coordinate first then the y-coordinate (i.e., (x,y)).
.
In your problem, you are given two ordered pairs -- the first is missing the y-coordinate and the second is missing the x-coordinate:
ordered pairs(0, )
and
(3, )
.
To find the missing coordinate, substitute the provided point into the given equation and solve for the missing point.
.
The first pair:(0, )
Here, they give you x is equal to zero. Substitute it in and solve for y.
y = 2x + 7
y = 2(0) + 7
y = 0 + 7
y = 7
First ordered pair is (0,7)
.
The second pair:(3, )
Here, they give you x is equal to three. Substitute it in and solve for y.
y = 2x + 7
y = 2(3) + 7
y = 6 + 7
y = 13
Second ordered pair is (3, 13)